SPECTROSCOPIC IDENTIFICATION OF NEW BENZYL-TYPE RADICALS IN CORONA EXCITATION: CHLORO-α-METHYLBENZYL RADICALS

YOUNG WOOK YOON, SEUNG WOON LEE, \underline{SANG KUK LEE}, Department of Chemistry, Pusan National University, Pusan 609-735, South Korea; and GI WOO LEE, Korea Basic Science Institute, Pusan 609-735, Korea.

Using a pinhole-type glass nozzle coupled with corona excited supersonic expansion (CESE), precursors o- and p-chloro-ethylbenzenes seeded in a large amount of inert carrier gas helium were corona discharged to produce jet-cooled benzyltype radicals. The vibronic emission spectra were recorded with a long path monochromator to observe vibronic bands of benzyl-type radicals in the visible region. The analysis of the spectra shows the evidence of the o- and p-chlorobenzyl radicals as typical products in corona excitation as well as the o- and p-chloro-α-methylbenzyl radicals as unexpected species.^{a} After subtracting the bands belonging to chlorobenzyl radicals, we could construct the vibronic emission spectra of the o- and p-chloro-α-methylbenzyl radical in the D_{1} → D_{0} electronic transition at the gas phase. By comparing with those of ab initio calculation and of the known vibrational mode frequencies with similar structures, we could obtain the electronic transition and assign the vibrational modes to the bands observed.

aH. G. Ahn, G. W. Lee, and S. K. Lee J. Phys. Chem. A \underline{112} (51), 13427 (2008).